OpenStudy (anonymous):

Mike was working on solving the exponential equation 37^x = 12; however, he is not quite sure where to start. Using complete sentences, describe to Mike how to solve this equation and how solving would be different if the bases were equal.

OpenStudy (anonymous):

12 does not have a base, but to solve \[37^x=12\] in one step you can write one of two things 1) \(x=\log_{37}(12)\) which actually tells you nothing, just restates the problem or 2) \(x=\frac{\log(12)}{\log(37)}\) by the change of base formula

OpenStudy (anonymous):

how do you do #2?

OpenStudy (anonymous):

not sure what you mean "how do you do" if you want a decimal for the second one you use a calculator, or wolfram hold on i will link

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=log%2812%29%2Flog%2837%29 looks like about \(.688\)

OpenStudy (anonymous):

I mean can you show me how do the second choice..

OpenStudy (anonymous):

Can you show me how to do both equations ?

OpenStudy (anonymous):

you mean how i got from \[37^x=12\] to \[x=\frac{\log(12)}{\log(37)}\] or how i computed that number?

OpenStudy (anonymous):

actually let me answer both questions first is it always the case that if \[b^x=A\] then \[x=\frac{\log(A)}{\log(b)}\] it is the log of the total divided by the log of the base that is how i went from \[37^x=12\] directly to \[x=\frac{\log(12)}{\log(37)}\]

OpenStudy (anonymous):

both

OpenStudy (anonymous):

the question of how you get a number out of this is by using a calculator

OpenStudy (anonymous):

for the first one turning \[37^x=12\] in to \[\log_{37}(12)=x\] is because \[b^x=A\iff \log_b(A)=x\]

OpenStudy (anonymous):

you should be able so switch between those forms easily for example \[2^x=8\iff \log_2(8)=3\] and \[10^{-3}=.001\iff\log_{10}(.001)=-3\]

OpenStudy (anonymous):

I am so confused... I have to put this in complete sentences and I don't understand what you're doing therefore I can't put it in complete sentences..

OpenStudy (anonymous):

lets go real slow

OpenStudy (anonymous):

the first thing you have is \(37^x=12\) right ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok now you have no idea what number to raise 37 to to get 12,

OpenStudy (anonymous):

we can simply rewrite the statement \[37^x=12\] as an equivalent statement \[\log_{37}(12)=x\] they say exactly the same thing and you don't know either of them

OpenStudy (anonymous):

before we go on, lets do one we do know we know \[2^4=16\] right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

that means \[\log_2(16)=4\] they say exactly the same thing, they just look different

OpenStudy (anonymous):

also we know \[10^{-3}=\frac{1}{10^3}=0.001\] that means \[\log_{10}(.001)=-3\] again they say the same thing, they just look different

OpenStudy (anonymous):

unfortunately \[x=\log_{37}(12)\] which is the same as \[37^x=12\] is not a number you know, so you will not be able to compute this number without a calculator on your calculator you have log base ten not log base 37

OpenStudy (anonymous):

ohk.. I think I understand.

OpenStudy (anonymous):

so what are you going to do to come up with \(\log_{37}(12)\)? you have to use the "change of base formula" and put \[x=\frac{\log(12)}{\log(37)}\]which you can compute using a calculator

OpenStudy (anonymous):

.6881